# MemoryBench: A Benchmark for Memory and Continual Learning in LLM Systems

### v2

@@ 3.2 Experimental Setup
Evaluation Metrics. MemoryBench follows the original evaluation metrics of each dataset. However, because certain datasets (JuDGE, IdeaBench, and SciTechNews) have multiple evaluation criteria in their official settings, we simplify the process by employing a LLM-as-Judge paradigm to merge these metrics into a single ranging from 1 to 10. We present the prompts used for this aggregation in A.1. For the overall evaluation of a task or domain, we apply min-max normalization or compute the z score to all results within each dataset before calculating the average performance of all test cases.
On-policy and Off-policy Settings. In each experiment, we first feed all the initial task context (i.e., the corpus for each training case) to a LLMsys, and then explore two settings to evaluate it. The first one is off-policy learning. We first run the backbone LLM of the LLMsys on all training cases and simulate 1 dialog session per each case to form the feedback logs. At each time step, we feed one batch of training cases and corresponding feedback sessions (one per each case) to the LLMsys. The second one is an on-policy learning setting. In each time step, we feed a batch of randomly sampled training cases to the LLMsys and collect feedback. The feedback logs are then fed to the LLMsys so that it can update its memory or parameters before the next time step. However, as discussed in Section 3.3 and A.3.8, some memory-based LLMsys are too slow that running on-policy experiments for them on all datasets would take days or even weeks. Therefore, we only reported the on-policy results for LLMsys that can finish the tasks within reasonable time (i.e., less than a day) in A.3. In MemoryBench, we simulate user feedback directly with a LLM-as-user paradigm using a strong LLM (i.e., Qwen-32B in this paper) and the evaluation metadata of each task. More details can be found in A.2 and A.3.
@@ 3.3 Results and Discussions
Usefulness of simulated user feedback.
Figure 2 presents the overall performance results (in off-policy settings, more detailed and on-policy results can be found in A.3) of all the baselines tested in our experiments on different partitions of MemoryBench. Note that the results for Mem0 is missing on Open-Domain and LiSo because the context lengths of those task data are so long that Mem0 cannot process in its memory and produce responses in reasonable time (more details can be found in A.3.8). While the best performing methods may vary across different domains and task formats, LLMsys with memory support mostly outperformed the Vanilla method that uses the backbone LLM without memory. To further validate that the reliability of the simulated user feedback, we conducted experiments comparing the performance of the backbone LLM (i.e., Vanilla) on each task case with and without feedback (results can be found in A.3.5). Our experiments show that the LLM-as-user paradigm could produce reasonably accurate feedback to LLMsys’s responses that are indeed useful for LLMsys to improve their performance. Yet, as shown in Figure 2, on SiLO, the performance of Vanilla is better than all memory-based LLMsys, which indicates that the effectiveness of existing LLMsys utilizing user feedback could vary significantly across tasks with different formats.
@@
Despite of the rising popularity of memory-based LLMsys, our experiments on MemoryBench show that the SOTA memory-based LLMsys are still unsatisfying. As shown in Figure 2, none of the advanced memory-based LLMsys (i.e., A-Mem, Mem0, or MemoryOS) can consistently outperform RAG baselines that simply use all task context and feedback logs as retrieval corpus. To understand whether this phenomenon is related to the types of feedback signals we collected, we also conducted experiments with the simulated action feedback such as “like” and “copy”, where we observed similar results (more details in A.3.9). This is contradicting to the good results reported by former studies (Xu et al., 2025a; Chhikara et al., 2025; Kang et al., 2025). After a careful examination of our experimental design and previous studies, we find several differences in our experimental design. First, previous studies only tested their system on reading comprehension tasks that involve long input context and require short answers (e.g., Locomo). If we only look at their performance on Locomo in our experiments, they indeed outperformed naive RAG baselines with BM25 (see Table 9 in A.3). However, Locomo is only one of the datasets tested in MemoryBench and many datasets have different input and output formats with it. This indicates that the generalizability of these SOTA memory-based LLMsys is limited. Second, MemoryBench provides both the task context and feedback logs from previous tasks to LLMsys. Existing memory-based LLMsys do not differentiate the historical task logs with current task context, which makes the use of feedback logs insufficient and even leads to additional noise to the system. While RAG baselines do not have sophisticated mechanism to analyze and utilize different types of task context and user feedback, their retrieval systems (i.e., term-based index systems or vector retrieval engines) are strong in handling large-scale data collection and retrieving relevant information efficiently. Thus, their performance are comparable to many advanced memory methods on MemoryBench.
@@
Figure 3: Time consumption of different LLMsys in off-policy experiments on four task-format-base partitions of MemoryBench. The bars show the average time per case for memory operations (Memory Time, in blue) and prediction on the test set (Predict Time, in orange), measured in seconds. Experiments were done with $\timesA8$00-80G GPU (details in A.3.8).
@@
In our experiments, we notice that both the effectiveness and the efficiency of existing memory-based LLMsys are far from perfect, especially for continual learning with increasing amount of user feedback. First, we are surprised to find that the time of existing memory-based methods used to memorize input context and conducting online inference are extremely long and varied inconsistently across tasks. Figure 3 shows the Memory Time (i.e., the time a LLMsys needs to process the behavior logs and the task context of one test case) and the Predict Time (i.e., the time a LLMsys needs to generate a response to a test case after the memory process) of different LLMsys on tasks with different formats. As shown in the figure, Mem0 has small memory time on tasks with long inputs (i.e., LiSo, LiLo), but longer time on tasks with short inputs (i.e., SiLo, SiSo). And comparing to other baselines, its inference speed is abnormally large on LiLo. MemoryOS exhibited better consistency in terms of inference time, but it took much longer time in memory construction (mostly longer than 17 seconds per case). A-MEM is the most efficient memory-based LLMsys in our experiments, probably thanks to its simplified memory construction and retrieval process. Yet, its effectiveness is not better than naive RAG baselines in many cases. Also, in our experiments, the performance of LLMsys fluctuated significantly during the training process, especially on Academic and Legal domains (see in A.3.7). This means that existing methods cannot effectively filter and utilize user feedback in continual learning. One possible reason is that, compared to Open domain, feedback logs on vertical domains might be more difficult to interpret as it requires more domain knowledge. Overall, our experiments on MemoryBench show that the memory mechanisms of existing LLMsys are not capable of handling multiple types of declarative and procedural memory at the same time.
@@ 4 Related Work
LLM Memory. The studies of LLM memory mostly focus on constructing external information management systems that help LLMsys accumulate knowledge, process historical experience, make informed decisions, etc. (Zhang et al., 2025d). A typical LLM memory system support multiple operations such as writing(Zhong et al., 2024), summarizing (Liu et al., 2023), updating (Xu et al., 2025a), deleting (Park et al., 2023), and retrieval (Packer et al., 2023). The applications of memory cover almost all use cases of LLMsys and agents, including social simulation (Zhang et al., 2025b), personal assistant (Lu et al., 2023), gaming (Yan et al., 2023), and much more.
Continual Learning for LLM. The concept of continual learning, especially those based on user feedback, is old to AI but relatively new to LLM. The majority of studies in LLM literature focuses on the pre-training Devlin et al. (2019) and post-training (Achiam et al., 2023) of LLMs based on SFT (Hu et al., 2022), RL (Guo et al., 2025a) or other advanced algorithms. There are hundreds of works on continual learning algorithms that adapt LLM to unseen data distributions (Shi et al., 2024), but the number of papers on continually optimizing LLMsys with user feedback is limited. One possible reason is that no commercial LLM service company has released their user logs, and there isn’t a comprehensive benchmark that has user feedback for LLMsys.
Feedback Simulation. User feedback simluation has been widely studied in IR literature for more than fifty years (Balog & Zhai, 2023), particularly on evaluation (Zhang et al., 2017), information seeking activities (Zhang et al., 2025a), search and recommendation behaviors (Craswell et al., 2008), conversational agents (Wang et al., 2024), etc. Traditional user simulation frameworks are often designed for well-developed yet simple user interfaces such as search engine and recommendation ranked lists (Wang et al., 2013). As the UI designs of LLMsys gradually become homogeneous, there are more and more studies on user behavior analysis for LLMsys (Liu et al., 2024). With the help of advanced human simulation (Wang et al., ) and LLM-as-judge frameworks (Li et al., 2024b), it is now possible to simulate user feedback to LLMsys with reasonably good quality.
@@ 7 Reproducibility Statement
To ensure reproducibility, all data and pre-processing scripts used in MemoryBench are open-sourced in https://huggingface.co/datasets/THUIR/MemoryBench and https://github.com/LittleDinoC/MemoryBench. Experiment details can be found in the appendix, and we have also release our implementation of the baselines in the github repository. We hope these measures can facilitate the replication of our results and foster future studies on memory and continual learning for LLMsys.
@@ A.1 Data Preparation and Evaluation Protocols
Table 2 reports the statistical information of the datasets included in MemoryBench. To measure text length, we compute the number of tokens after tokenized by the Qwen3-8B tokenizer. For inputs, we combine the query with the corpus (if available), while for outputs, we use the length of the golden answers. For datasets without golden answers (e.g., HelloBench, IdeaBench, WritingBench), we record the average length that Qwen3-8B actually generates. We set 600 tokens as the boundary for distinguishing between short and long inputs/outputs. In the LLM-as-Judge evaluation, we follow the official prompt templates provided by each dataset. We use the officially released WritingBench-Critic-Model-Qwen-7B [^3] https://huggingface.co/AQuarterMile/WritingBench-Critic-Model-Qwen-7B for WritingBench, and gpt-4o-mini for the other datasets.
@@ A.1.1 Details of Data Preparation
  • LexEval (Li et al., 2024a): LexEval is a comprehensive Chinese legal benchmark. We select three subsets labeled 5-1, 5-2, and 5-4, corresponding to Summary Generation, Judicial Analysis Generation, and Open-ended Question Answering. The evaluation metric is the Rouge-L score compared to the golden answers.
  • JuDGE (Su et al., 2025): JuDGE is a Chinese Legal Judgment Document Generation evaluation dataset. LLMsys need to generate a complete Judgment Document based on the facts of legal cases. We incorporate both the JuDGE training and test sets as a subset of MemoryBench. JuDGE uses a broad range of evaluation metrics, focusing primarily on differences from the golden documents annotated by human experts. These metrics include the similarity of the Reasoning and Judgment Sections (assessed via METEOR and BERTScore), the accuracy of legal references and charges (measured by recall, precision, and F1), and the relative differences in fines and prison sentences.
  • LimitGen-Syn (Xu et al., 2025b): LimitGen systematically evaluates how LLMs identify limitations in scientific papers, particularly in AI research. It includes a human-annotated subset (LimitGen-Human) and a synthetic subset (LimitGen-Syn). Because the former requires multiple dialogue sessions to generate answers, and to align with the other MemoryBench datasets, we only choose the LimitGen-Syn subset. We follow the original paper’s setup by specifying in the prompt which limitation subtype (e.g., Methodology, Experimental Design) should be recognized, and we instruct LLMsys to generate three limitations in JSON format. Following the original paper, we use an LLM-based rating metric (ranging from 0–5) during evaluation. We first match the limitation subtype at a coarse-grained level, then compare these matches to the ground truth to assign scores, and finally select the highest scoring limitation among the three generated by LLMsys as the final result.
  • WritingBench (Wu et al., 2025d): WritingBench is a multi-domain benchmark that evaluates writing proficiency in both English and Chinese using dynamic evaluation rules, with the critic model required to produce scores from 1 to 10. We build three MemoryBench subsets from WritingBench:
@@
  • WritingPrompts (Fan et al., 2018): WritingPrompts is a story generation dataset, from which we use the first 2000 items in the test set. Evaluation relies on the METEOR score (Banerjee & Lavie, 2005), comparing generated stories to human-written ones.
  • Locomo (Maharana et al., 2024): Locomo is a dataset designed to evaluate the long-term memory capability of LLMsys. It includes 10 long dialogues, each containing around 200 questions. We treat each dialogue (session-level) as the corpus and each question as a user query. We then calculate the Average F1 score of all answers against the gold answers (denoted in the original paper as “Answer Prediction (F1) – Overall”).
  • DialSim (Kim et al., 2025): DialSim aims to evaluate the Long-Term Multi-Party Dialogue Understanding of LLMsys. We adopt the official v1.1 version of the dataset, and divide its content by the TV show source into three subsets: friends, bigbang, and theoffice. Each subset’s corpus comprises the complete scripts of the corresponding TV show, and the user queries are drawn from Fan Quiz-type questions. We only consider questions that are answerable and with time to ensure alignment between the corpus and the queries. In the official DialSim prompt, the full dialogue history is given directly to the model; however, within MemoryBench, we require each baseline to handle memory automatically. To accommodate this, we minimally modify the original prompt by removing the dialogue history section and leaving everything else unchanged, as follows:
@@
  • SciTechNews (Cardenas et al., 2023): SciTechNews is a dataset of Science Journalism for the General Audience. We use its test set and follow the evaluation settings proposed in the original SciTechNews paper and the JRE-L framework  (Jiang et al., 2025). Our metrics include two primary dimensions. The first examines how closely the generated journalism aligns with the human-provided golden answers, using Rouge-L and BERTScore-F1  (Zhang et al., 2020) to calculate factual accuracy. The second relies on readability measures, such as CLI  (Coleman & Liau, 1975), FKGL  (Kincaid et al., 1975), and DCRS  Dale & Chall (1948), to assess the readability of the generated journalism. A lower readability score indicates that the article is better suited for science news.
  • IdeaBench (Guo et al., 2025b): IdeaBench is a benchmark that evaluates the generation of research ideas by LLMsys, requiring LLMsys to create new research hypotheses or ideas based on a set of reference abstracts. Its evaluation framework is multifaceted and mainly examines how the generated ideas align with the “real” ideas in the target paper under various quality metrics. The primary metric is the “Insight Score,” obtained through a two-step procedure: first, LLM-as-Judge ranks all ideas against user-specified criteria (such as “novelty” and “feasibility”); second, the final score is determined by the relative ranking of the target paper’s ideas. Additional evaluations include semantic similarity (scored by BERTScore F1) and idea overlap (also scored by LLM-as-Judge) as reference metrics.
  • HelloBench (Que et al., 2024): HelloBench evaluates the capability of LLM to generate long-form text across five tasks: open-ended question answering, chat, summarization, text completion, and heuristic text generation. Following the original paper’s recommendation, we employ the “Average evaluation results of Checklists” method, which correlates most strongly with human judgments. In this approach, the LLM assigns various dimension scores based on a checklist for each question, and the average of these scores serves as the final result. From HelloBench, we derive three MemoryBench subsets:
@@
  • NF-Cats (Bolotova et al., 2022): NF-Cats is a Non-Factoid Question-Answering dataset. We use its test set and follow the LLM as Judge evaluation approach recommended by NTCIR-18 (Chen et al., 2025). The prompt template for this evaluation is as follows:

### v4

@@ 3.2 Experimental Setup
Evaluation Metrics. MemoryBench follows the original evaluation metrics of each dataset. Many of these datasets (e.g., JuDGE, LexEval, IdeaBench, SciTechNews) have human-crafted ground truth and official evaluation protocols that are used for comparison. These gold standards often represent near-human-optimal responses and provide important information about the performance ceiling for human and LLMsys on each task (please refer to the dataset references for details). Because many datasets in MemoryBench have multiple evaluation criteria in their official settings, for simplicity and better visualization, we employ a LLM-as-Judge paradigm to merge these metrics into a single ranging from 1 to 10. We present the prompts used for this aggregation in A.1. Then, for the overall evaluation of a task or domain, we apply min-max normalization or compute the z score to all results within each dataset before calculating the average performance of all test cases. We also report the original metric performance on each dataset in our off-policy experiments in Table 12, A.3.
On-policy and Off-policy Settings. In each experiment, we first feed all the initial task context (i.e., the corpus for each training case) to a LLMsys, and then explore two settings to evaluate it. The first one is off-policy learning. We first run the backbone LLM of the LLMsys on all training cases and simulate 1 dialog session per each case to form the feedback logs. At each time step, we feed one batch of training cases and corresponding feedback sessions (one per each case) to the LLMsys. The second one is an on-policy learning setting. In each time step, we feed a batch of randomly sampled training cases to the LLMsys and collect feedback. The feedback logs are then fed to the LLMsys so that it can update its memory or parameters before the next time step. However, as discussed in Section 3.3 and A.3.8, some memory-based LLMsys are too slow that running on-policy experiments for them on all datasets would take days or even weeks. Therefore, we only reported the on-policy results for LLMsys that can finish the tasks within reasonable time (i.e., less than a day) in A.3. In MemoryBench, we simulate user feedback directly with a LLM-as-user paradigm using a strong LLM (i.e., Qwen-32B in the main results of this paper if not explicitly mentioned otherwise, but we also provide the feedback by Mistral3.2-24B in the data link [^3] https://huggingface.co/datasets/THUIR/MemoryBench) and the evaluation metadata of each task. More details can be found in A.2 and A.3.
For simplicity, if not explicitly stated otherwise, all experiments reported in the main content of this paper are conducted in the off-policy setting with explicit verbose feedback only. We also provide more experiment results on on-policy settings and other types of feedback (e.g., explicit actions and implicit actions such as clicks on “like” and “copy” buttons) in A.3.
@@ 3.3 Results and Discussions
Usefulness of simulated user feedback.
Figure 3 presents the overall performance results (off-policy settings, explicit verbose feedback only, more detailed and on-policy results can be found in A.3) of all the baselines tested in our experiments on different partitions of MemoryBench. Note that the results for Mem0 is missing on Open-Domain and LiSo because the context lengths of those task data are so long that Mem0 cannot process in its memory and produce responses in reasonable time (more details can be found in A.3.8). While the best performing methods may vary across different domains and task formats, LLMsys with memory support mostly outperformed the Vanilla method that uses the backbone LLM without memory. To further validate that the reliability of the simulated user feedback, we conducted experiments comparing the performance of the backbone LLM (i.e., Vanilla) on each task case with and without feedback (results can be found in A.3.5). Our experiments show that the LLM-as-user paradigm could produce reasonably accurate feedback to LLMsys’s responses that are indeed useful for LLMsys to improve their performance. Yet, as shown in Figure 3, on SiLO, the performance of Vanilla is better than all memory-based LLMsys, which indicates that the effectiveness of existing LLMsys utilizing user feedback could vary significantly across tasks with different formats.
@@
Despite of the rising popularity of memory-based LLMsys, our experiments on MemoryBench show that the SOTA memory-based LLMsys are still unsatisfying. As shown in Figure 3, none of the advanced memory-based LLMsys (i.e., A-Mem, Mem0, or MemoryOS) can consistently outperform RAG baselines that simply use all task context and feedback logs as retrieval corpus. To understand whether this phenomenon is related to the types of feedback signals we collected, we also conducted experiments with the simulated action feedback such as “like” and “copy”, where we observed similar results (more details in A.3.9). This is contradicting to the good results reported by former studies (xu2025mem; chhikara2025mem0; kang2025memory). After a careful examination of our experimental design and previous studies, we find several differences in our experimental design. First, previous studies only tested their system on reading comprehension tasks that involve long input context and require short answers (e.g., Locomo). They manually tailored their methods to handle such tasks using special chunking, summarization, and hierarchical memory-clustering techniques to handle long input context. In contrast, the RAG baselines used in our paper simply build their retrieval corpus by splitting the context by dialogs or messages without any task-specific design. If we only look at the models’ performance on Locomo in our experiments, existing memory-based methods indeed outperformed naive RAG baselines with BM25 (see Table 12 in A.3). However, Locomo is only one of the datasets tested in MemoryBench and many datasets have different input and output formats with it. This indicates that the generalizability of these SOTA memory-based LLMsys is limited. Second, MemoryBench provides both the current task context and the feedback logs from previous tasks to LLMsys. This makes it significantly different from previous benchmarks where only the current task context is provided. As discussed in Section 2.1, the task context provided by previous benchmarks are usually factual information independent to how the system performs with respect to the task in the history (i.e., declarative memory). The feedback logs provided in MemoryBench are non-factual information describing how the system performed and what it has done well or badly in historical tasks (i.e., procedural memory). Existing memory-based LLM systems such as Mem0 and MemoryOS simply treat all inputs as declarative memory and develop memory mechanisms accordingly. This limits their ability to understand and utilize procedural memory, i.e., the feedback logs in MemoryBench. While RAG baselines do not have sophisticated mechanism to analyze and utilize different types of task context and user feedback, their retrieval systems (i.e., term-based index systems or vector retrieval engines) are strong in handling large-scale data collection and retrieving relevant information efficiently. Thus, their performance are comparable to many advanced memory methods on MemoryBench.
@@
Figure 4: Time consumption of different LLMsys in off-policy experiments on four task-format-base partitions of MemoryBench. The bars show the average time per case for memory operations (Memory Time, in blue) and prediction on the test set (Predict Time, in orange), measured in seconds. Note that the construction of memory indexes or retrieval indexes are also considered as a part of the Memory Time. Experiments were done with $\timesA8$00-80G GPU (details in A.3.8).
@@
In our experiments, we notice that both the effectiveness and the efficiency of existing memory-based LLMsys are far from perfect, especially for continual learning with increasing amount of user feedback. First, we are surprised to find that the time of existing memory-based methods used to memorize input context and conducting online inference are extremely long and varied inconsistently across tasks. Figure 4 shows the Memory Time (i.e., the time a LLMsys needs to process the behavior logs and the task context of one test case, including the time of index construction in RAG methods) and the Predict Time (i.e., the time a LLMsys needs to retrieve memory and generate a response to a test case) of different LLMsys on tasks with different formats. As shown in the figure, Mem0 has small memory time on tasks with long inputs (i.e., LiSo, LiLo), but longer time on tasks with short inputs (i.e., SiLo, SiSo). And comparing to other baselines, its inference speed is abnormally large on LiLo. MemoryOS exhibited better consistency in terms of inference time, but it took much longer time in memory construction (mostly longer than 17 seconds per case). A-MEM is the most efficient memory-based LLMsys in our experiments, probably thanks to its simplified memory construction and retrieval process. Yet, its effectiveness is not better than naive RAG baselines in many cases. Also, in our experiments, the performance of LLMsys fluctuated significantly during the training process, especially on Academic and Legal domains (see in A.3.7). This means that existing methods cannot effectively filter and utilize user feedback in continual learning. One possible reason is that, compared to Open domain, feedback logs on vertical domains might be more difficult to interpret as it requires more domain knowledge. Overall, our experiments on MemoryBench show that the memory mechanisms of existing LLMsys are not capable of handling multiple types of declarative and procedural memory at the same time.
@@ 4 Related Work
Memory-based LLM systems. The studies of LLM memory mostly focus on constructing external information management systems that help LLMsys accumulate knowledge, process historical experience, make informed decisions, etc. (ZhangSurvey). In general, advanced memory-based LLMsys consists of multiple modules for memory processing, including but not limited to a memory processor that take user-system conversations as input to build memory entries/documents; a memory database that store and indexes memory entries; a memory retriever to retrieve relevant memory entries given a specific context; and a generator that takes the retrieved memory and current task context to generate the final responses. A typical LLM memory system support multiple operations such as writing(zhong2024memorybank), summarizing (liu2023think), updating (xu2025mem), deleting (park2023generative), and retrieval (packer2023memgpt). It operates its memory database throughout its interactions with the environments, and utilize relevant memory retrieved from the database to better improve its responses to the environment inputs. The applications of memory cover almost all use cases of LLMsys and agents, including social simulation (zhang2025chinesecourtsimulationllmbased), personal assistant (lu2023memochat), gaming (yan2023larp), and much more.
Continual Learning for LLM. The concept of continual learning, especially those based on user feedback, is old to AI but relatively new to LLM. The majority of studies in LLM literature focuses on the pre-training devlin2019bert and post-training (achiam2023gpt) of LLMs based on SFT (hu2022lora), RL (guo2025deepseek) or other advanced algorithms. There are hundreds of works on continual learning algorithms that adapt LLM to unseen data distributions (shi2024continual), but the number of papers on continually optimizing LLMsys with user feedback is limited. One possible reason is that no commercial LLM service company has released their user logs, and there isn’t a comprehensive benchmark that has user feedback for LLMsys. While there has been studies analyzing LLMsys’s ability in multi-turn conversations that involve utterances that could serve as user feedback (you2024llm), they considered these utterances only as local context and discard them after the system finish each task. To the best of our knowledge, MemoryBench is the first benchmark that can evaluate the continual learning ability of LLMsys on user feedback logs.
Feedback Simulation. User feedback simluation has been widely studied in IR literature for more than fifty years (krisztain2023simulation), particularly on evaluation (zhang2017information), information seeking activities (zhang2025exploring), search and recommendation behaviors (craswell2008experimental), conversational agents (wang2024depth), etc. Traditional user simulation frameworks are often designed for well-developed yet simple user interfaces such as search engine and recommendation ranked lists (wang2013incorporating). As the UI designs of LLMsys gradually become homogeneous, there are more and more studies on user behavior analysis for LLMsys (Liu2024chatgptbehavior). With the help of advanced human simulation (wangsurvey) and LLM-as-judge frameworks (li2024llmsasjudgescomprehensivesurveyllmbased), it is now possible to simulate user feedback to LLMsys with reasonably good quality.
@@ 7 Reproducibility Statement
To ensure reproducibility, all data and pre-processing scripts used in MemoryBench are open-sourced in https://anonymous.4open.science/r/MemoryBench-Dataset-BF1F and https://anonymous.4open.science/r/MemoryBench/. Experiment details can be found in the appendix, and we have also release our implementation of the baselines in the github repository. We hope these measures can facilitate the replication of our results and foster future studies on memory and continual learning for LLMsys.
@@ A.1 Data Preparation and Evaluation Protocols
Table 2 reports the statistical information of the datasets included in MemoryBench. To measure text length, we compute the number of tokens after tokenized by the Qwen3-8B tokenizer. For inputs, we combine the query with the corpus (if available), while for outputs, we use the length of the golden answers. For datasets without golden answers (e.g., HelloBench, IdeaBench, WritingBench), we record the average length that Qwen3-8B actually generates. We set 600 tokens as the boundary for distinguishing between short and long inputs/outputs. In the LLM-as-Judge evaluation, we follow the official prompt templates provided by each dataset. We use the officially released WritingBench-Critic-Model-Qwen-7B [^4] https://huggingface.co/AQuarterMile/WritingBench-Critic-Model-Qwen-7B for WritingBench, and DeepSeek-V3 for the other datasets.
Table 3: The categorization of datasets in MemoryBench in terms of memory types. Declarative, Procedural, Semantic, Episodic are abbreviations for Declarative, Procedural, Semantic (user independent memory), and Episodic memory (user-specific memory).
Dataset Name Memory Type
Declarative Procedural
Semantic Episodic.
LexEval $\checkmark$ $\times$ $\times$
JUDGE $\checkmark$ $\times$ $\times$
LimitGen-Syn $\checkmark$ $\times$ $\times$
WritingBench $\checkmark$ $\times$ $\times$
WritingPrompts $\checkmark$ $\times$ $\times$
Locomo $\times$ $\checkmark$ $\times$
DialSim $\checkmark$ $\checkmark$ $\times$
SciTechNews $\checkmark$ $\times$ $\times$
IdeaBench $\checkmark$ $\times$ $\times$
HelloBench $\checkmark$ $\checkmark$ $\times$
NF-Cats $\checkmark$ $\times$ $\times$
@@ A.1.1 Details of Data Preparation
Here we introduce the detailed descriptions of each datasets used in MemoryBench. A categorization of the datasets based on the taxonomy of memory used in Section 2.1 is shown in Table 3. Specificially,
  • LexEval (NEURIPS2024_2cb40fc0-lexeval): LexEval is a comprehensive Chinese legal benchmark. We select three subsets labeled 5-1, 5-2, and 5-4, corresponding to Summary Generation, Judicial Analysis Generation, and Open-ended Question Answering. The evaluation metric is the Rouge-L score compared to the golden answers.
  • JuDGE (judge2025): JuDGE is a Chinese Legal Judgment Document Generation evaluation dataset. LLMsys need to generate a complete Judgment Document based on the facts of legal cases. We incorporate both the JuDGE training and test sets as a subset of MemoryBench. JuDGE uses a broad range of evaluation metrics, focusing primarily on differences from the golden documents annotated by human experts. These metrics include the similarity of the Reasoning and Judgment Sections (assessed via METEOR and BERTScore), the accuracy of legal references and charges (measured by recall, precision, and F1), and the relative differences in fines and prison sentences.
  • LimitGen-Syn (xu-etal-2025-llms-identify-limitgen): LimitGen systematically evaluates how LLMs identify limitations in scientific papers, particularly in AI research. It includes a human-annotated subset (LimitGen-Human) and a synthetic subset (LimitGen-Syn). Because the former requires multiple dialogue sessions to generate answers, and to align with the other MemoryBench datasets, we only choose the LimitGen-Syn subset. We follow the original paper’s setup by specifying in the prompt which limitation subtype (e.g., Methodology, Experimental Design) should be recognized, and we instruct LLMsys to generate three limitations in JSON format. Following the original paper, we use an LLM-based rating metric (ranging from 0–5) during evaluation. We first match the limitation subtype at a coarse-grained level, then compare these matches to the ground truth to assign scores, and finally select the highest scoring limitation among the three generated by LLMsys as the final result.
  • WritingBench (wu2025writingbench): WritingBench is a multi-domain benchmark that evaluates writing proficiency in both English and Chinese using dynamic evaluation rules, with the critic model required to produce scores from 1 to 10. We build three MemoryBench subsets from WritingBench:
@@
  • WritingPrompts (fan-etal-2018-hierarchical-writingprompts): WritingPrompts is a story generation dataset, from which we use the first 2000 items in the test set. Evaluation relies on the METEOR score (banerjee-lavie-2005-meteor), comparing generated stories to human-written ones.
  • Locomo (maharana2024evaluatinglongtermconversationalmemory): Locomo is a dataset designed to evaluate the long-term memory capability of LLMsys. It includes 10 long dialogues, each containing around 200 questions. We treat each dialogue (session-level) as the corpus and each question as a user query. We then calculate the Average F1 score of all answers against the gold answers (denoted in the original paper as “Answer Prediction (F1) – Overall”).
  • DialSim (kim2025dialsimrealtimesimulatorevaluating): DialSim aims to evaluate the Long-Term Multi-Party Dialogue Understanding of LLMsys. We adopt the official v1.1 version of the dataset, and divide its content by the TV show source into three subsets: friends, bigbang, and theoffice. Each subset’s corpus comprises the complete scripts of the corresponding TV show, and the user queries are drawn from Fan Quiz-type questions. We only consider questions that are answerable and with time to ensure alignment between the corpus and the queries. In the official DialSim prompt, the full dialogue history is given directly to the model; however, within MemoryBench, we require each baseline to handle memory automatically. To accommodate this, we minimally modify the original prompt by removing the dialogue history section and leaving everything else unchanged, as follows:
@@
  • SciTechNews (cardenas-etal-2023-dont-scitechnews): SciTechNews is a dataset of Science Journalism for the General Audience. We use its test set and follow the evaluation settings proposed in the original SciTechNews paper and the JRE-L framework  (jiang-etal-2025-jre). Our metrics include two primary dimensions. The first examines how closely the generated journalism aligns with the human-provided golden answers, using Rouge-L and BERTScore-F1  (zhang2020bertscoreevaluatingtextgeneration) to calculate factual accuracy. The second relies on readability measures, such as CLI  (coleman1975computer), FKGL  (kincaid1975derivation), and DCRS  dale1948formula, to assess the readability of the generated journalism. A lower readability score indicates that the article is better suited for science news.
  • IdeaBench (ideabench): IdeaBench is a benchmark that evaluates the generation of research ideas by LLMsys, requiring LLMsys to create new research hypotheses or ideas based on a set of reference abstracts. Its evaluation framework is multifaceted and mainly examines how the generated ideas align with the “real” ideas in the target paper under various quality metrics. The primary metric is the “Insight Score,” obtained through a two-step procedure: first, LLM-as-Judge ranks all ideas against user-specified criteria (such as “novelty” and “feasibility”); second, the final score is determined by the relative ranking of the target paper’s ideas. Additional evaluations include semantic similarity (scored by BERTScore F1) and idea overlap (also scored by LLM-as-Judge) as reference metrics.
  • HelloBench (que2024hellobenchevaluatinglongtext): HelloBench evaluates the capability of LLM to generate long-form text across five tasks: open-ended question answering, chat, summarization, text completion, and heuristic text generation. Following the original paper’s recommendation, we employ the “Average evaluation results of Checklists” method, which correlates most strongly with human judgments. In this approach, the LLM assigns various dimension scores based on a checklist for each question, and the average of these scores serves as the final result. From HelloBench, we derive three MemoryBench subsets:
@@
  • NF-Cats (nfcats): NF-Cats is a Non-Factoid Question-Answering dataset. We use its test set and follow the LLM as Judge evaluation approach recommended by NTCIR-18 (chen2025overviewntcir18automaticevaluation). The prompt template for this evaluation is as follows: